OpenStudy (anonymous):

PLEASE HELP log33 + log3x = 5 ??

OpenStudy (whpalmer4):

Is that \[\log_{3}3+\log_{3}x=5\]?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Adding logs of two numbers is the same as taking the log of the product so we can rewrite that as \[\log(3x)=5\] Do you know how to solve that?

OpenStudy (whpalmer4):

We could also notice that log base 3 of 3 =1 because 3^1=3

OpenStudy (anonymous):

No, I'm terrible at these :/

OpenStudy (whpalmer4):

If we go that route, \[\log_{3}x=4\]

OpenStudy (anonymous):

lol from conic sections to logs in one day. what is next? riemann sums? baye's formula?

OpenStudy (anonymous):

that's not helping -___- haha

OpenStudy (whpalmer4):

Let's go that latter route. The log base 3 of x means find the number where 3^number gives you x. We know the number is 4, so \(3^4=x\)

OpenStudy (whpalmer4):

\[x=3^4=3*3*3*3=\]

OpenStudy (anonymous):

So the final answer would be 81 @whpalmer4

OpenStudy (whpalmer4):

Yep!

OpenStudy (anonymous):

Thanks!

OpenStudy (whpalmer4):

We'll do a more indepth log tutorial sometime, just not while I'm walking around in Costco :-)

OpenStudy (anonymous):

oooh get me some kirkland coffee !

OpenStudy (whpalmer4):

Do you buy the small pallet load, or the big one? :-)

OpenStudy (anonymous):

and those cheap san marzano tomatoes too

OpenStudy (whpalmer4):

Which location are you going to that has the tomatoes? Never seen them in CA